A thin uniform rod has a length of 0.520 m and is rotating in a circle on a frictionless table. The axis of rotation is perpendicular to the length of the rod at one end and is stationary. The rod has an angular velocity of 0.31 rad/s and a moment of inertia about the axis of 3.10×10^(−3) kg⋅m2 . A bug initially standing on the rod at the axis of rotation decides to crawl out to the other end of the rod. When the bug has reached the end of the rod and sits there, its tangential speed is 0.129 m/s . The bug can be treated as a point mass.

A. What is the mass of the rod?

B. What is the mass of the bug?
Answer #1

a)

moment of inertia of rod

I = (1/3) M L^2

3.1* 10^-3 = (1/3)* M * 0.52^2

M = 0.0344 kg

=====

b)

angular velocity

w = 0.129 / 0.52 = 0.2481 rad/s

now using conservation of angular momentum

3.1* 10^-3* 0.31 = (3.1* 10^-3 + m * 0.52^2)* 0.2481

m = 2.86* 10^-3 kg

=====

Comment before rate in case any doubt, will reply for sure.. goodluck